What is the potential at point A? Give your answer
Answer #1

Potential due to q1 = 9x is V1 = k*q1/r1 = k*9x/sqrt(9L^2+4L^2) = 9*k*x/(L*sqrt(13)) = 2.5*k*x/L

potential due to q2 is V2 = k*q2/r2= -k*3x/3L = -k*x/L

potential due to q3 is V3 = k*q3/r3 = k*2x/2L = k*x/L

potential due to q4 is V4 = k*q4/r4 = k*(-7x)/(sqrt(8)*L) = -2.5*k*x/L

due to q5 is V5 = k*q5/r5 = k*4x/sqrt(L^2+4L^2) = 1.8*k*x/L


due to q6 is V6 = k*q6/r6 = k*(-2x)/2L = -k*x/L


due to q7 is V7 = k*q7/r7 = k*3x/L = 3*k*x/L

due to q8 is V8 = k*q8/r8 = k*(-7x)/(sqrt(8)*L) = -2.5*k*x/L

due to q9 is V9 = k*q9/r9 = k*(-4x)/(sqrt(5)*L) = -1.8*k*x/L

V = V1+V2+V3+V4+V5+V6+V7+V8+V9


V = kx/L[(2.5-1+1-2.5+1.8-1+3-2.5-1.8)]

V = -0.5*k*x/L